# Instability of sphalerons in $\phi^{4}$ models with a false vacuum

### v1

@@ 2. Sphaleron in a false vacuum
[…]
Case $\tilde{a}>0$: the vacua are interchanged, with $\phi=\tilde{b}$ being the false vacuum and $\phi=-\tilde{b}$ the true vacuum, where $\tilde{b}<0<\tilde{a}<\tilde{b}$. The potential (3) in this case supports a solitary wave on a bright background, with $\phi\to\tilde{b}$ at $x\to\pm\infty$. This wave describes a ground state lump when it is stationary. Using similar steps as in the previous case, we can shift the potential to get
$)V(\phi)=\alpha\phi^{2}\big{(}\tfrac{1}{4}\phi^{2}+(\tfrac{1}{3}\tilde{a}-\tilde{b})\phi+\tilde{b}(\tilde{b}-\tilde{a}$$\big{)}$ (6)
@@
The reparameterization of the potential (5) is achieved by expressing the pair of non-zero roots as
@@
$).b=2\sqrt{\tilde{b}(\tilde{b}-\tilde{a})},\quad a=\tfrac{1}{2}{\rm arctanh}\Big{(}\sqrt{1-\tilde{a}/\tilde{b}}\big{/}(1-\tfrac{1}{3}\tilde{a}/\tilde{b}$$\Big{)}$. (8)
@@
with $\tilde{b}>\tilde{a}>0$ listed in Ref. [7]. A plot is shown in Fig. 1.
Figure 1. Potential with a false vacuum at $\phi=0$: $a=$ 0.5, 0.8, 1.5
@@ 2.1. Ground states
Figure 2. Lump solution (19) in the false-vacuum potential for $a=$ 5.5, 2.5, 1.0, 0.5
@@ 3.1. Spectrum of the perturbation potential
These eigenfunctions have a single peak, at $x=0$, for $a\gtrsim 1.00$, and a double peak for $a\lesssim 1.00$.
@@
The eigenvalue equation (53) can be investigated numerically to look for other solutions, which would describe symmetric eigenfunctions. We find that it possesses one positive solution, which we denote as $\lambda_{1}$. Sturm’s oscillation theorem implies that the corresponding symmetric eigenfunction, $\eta_{1}(x)=\tilde{\eta}_{1}(y)$, has two nodes. It is called an internal vibrational (or shape) mode. Fig. 7 shows a plot of the eigenvalue and corresponding eigenfunction.
@@
Since the continuous spectrum starts at $\lambda=4$, what remains to be determined is if there is another positive eigenvalue $\lambda_{2}>\lambda_{1}$ corresponding to an eigenfunction that has three nodes. The determining condition for the eigenvalue is simply that $y^{\sqrt{1-\lambda/4}}(\sinh(a)^{2}y+1)^{3}H(y)$ must vanish at $y=1$. We find that it does have a solution, which is plotted in Fig. 8. The corresponding eigenfunction $\eta_{2}(x)=\tilde{\eta}_{2}(y)$ is also plotted. This represents an additional internal vibrational mode.
@@ 4. Approximate analytical formula for the unstable mode
$\eta_{-1}(x)\simeq{\rm sech}(x)^{8}\cosh(2x)^{4}\big{(}4e^{-2a}\sinh(x)^{2}+1\big{)}^{3}\big{(}1+24e^{-4a}\ln{\rm sech}(x)\big{)}/\big{(}2e^{-2a}\sinh(x)^{2}+1\big{)}^{4}$ (68)

### v2

@@ 2. Sphaleron in a false vacuum
$V(\phi)=\alpha\phi^{2}\big(\tfrac{1}{4}\phi^{2}+(\tfrac{1}{3}|\tilde{a}|-\tilde{b})\phi+\tilde{b}(\tilde{b}-|\tilde{a}|)\big),\quad\tilde{b}>-\tilde{a}>0$ (5)
@@
Case $\tilde{a}>0$: the vacua are interchanged, with $\phi=\tilde{b}$ being the false vacuum and $\phi=-\tilde{b}$ the true vacuum, where $\tilde{b}<0<\tilde{a}<\tilde{b}$. The potential (3) in this case supports a solitary wave on a bright background, $\phi\to\tilde{b}$ at $x\to\pm\infty$. This wave describes a ground state lump when it is stationary. Using similar steps as in the previous case, we can shift the potential to get
$V(\phi)=\alpha\phi^{2}\big(\tfrac{1}{4}\phi^{2}+(\tilde{b}-\tfrac{1}{3}\tilde{a})\phi+\tilde{b}(\tilde{b}-\tilde{a})\big),\quad\tilde{b}>\tilde{a}>0$ (6)
@@
The reparameterization of the potential (5) is achieved by expressing its two positive roots as
@@
$).b=2\sqrt{\tilde{b}(\tilde{b}-|\tilde{a}|)},\quad a=\tfrac{1}{2}{\rm arctanh}\Big(\sqrt{1-|\tilde{a}|/\tilde{b}}\big/\big(1-\tfrac{1}{3}|\tilde{a}|/\tilde{b}\big$$\Big)$. (8)
@@
with $\tilde{b}>\tilde{a}>0$ listed in Ref. [7]. A plot of family (9) is shown in Fig. 1.
Figure 1. Potential (9) with a false vacuum at $\phi=0$: $a=$ 0.5, 0.8, 1.5
@@ 2.1. Ground states
Figure 2. Lump solution (19) in the false-vacuum potential (9) for $a=$ 5.5, 2.5, 1.0, 0.5
@@ 3.1. Spectrum of the perturbation potential
These eigenfunctions have a single peak, at $x=0$, for $a\lesssim 1.00$, and a double peak for $a\gtrsim 1.00$. For $a\simeq 1.00$, when the eigenfunction has no convexity at $x=0$, the corresponding eigenvalue is $\lambda_{-1}\simeq 1.30$.
@@
The eigenvalue equation (53) can be investigated numerically to look for other solutions, which would describe symmetric eigenfunctions. We find that it possesses one positive solution, which we denote as $\lambda_{1}$. Sturm’s oscillation theorem implies that the corresponding symmetric eigenfunction, $\eta_{1}(x)=\tilde{\eta}_{1}(y)$, has two nodes. It is called an internal vibrational (or shape) mode. As it has a positive eigenvalue, this mode is linearly stable. Fig. 7 shows a plot of the eigenvalue and corresponding eigenfunction.
@@
Since the continuous spectrum starts at $\lambda=4$, what remains to be determined is if there is another positive eigenvalue $\lambda_{2}>\lambda_{1}$ corresponding to an eigenfunction that has three nodes. The determining condition for the eigenvalue is simply that $y^{\sqrt{1-\lambda/4}}(\sinh(a)^{2}y+1)^{3}H(y)$ must vanish at $y=1$. We find that it does have a solution, which is plotted in Fig. 8. The corresponding eigenfunction $\eta_{2}(x)=\tilde{\eta}_{2}(y)$ is also plotted. This represents an additional internal vibrational mode, which is linearly stable.
@@ 4. Approximate analytical formula for the unstable mode
$\eta_{-1}(x)\simeq\big(2-{\rm sech}(x)^{2}\big)^{4}\big(4e^{-2a}\sinh(x)^{2}+1\big)^{3}\big(1+24e^{-4a}\ln{\rm sech}(x)\big)/\big(2e^{-2a}\sinh(x)^{2}+1\big)^{4}$ (68)